OpenStudy (anonymous):

Celia’s long distance company charges $5.95 per month plus $0.06 per minute. If Celia has budgeted $30 for long distance, what is the maximum number of minutes she can call long distance per month? A.) 405 minutes B.) 375 minutes C.) 400 minutes D.) 420 minutes

OpenStudy (mathstudent55):

Let c = cost Let m = number of minutes c = 5.95 + 0.06m She has $30 to spend, so the cost must be less than or equal to $30 5.95 + 0.06m <= 30 Can you solve this inequality for m?

OpenStudy (campbell_st):

well you need to subtract the monthly charge form the cost 30 - 5.95 = when you have this answer divide it by 0.06... the call cost in dollars

OpenStudy (anonymous):

= 400

OpenStudy (mathstudent55):

Correct

OpenStudy (campbell_st):

yep

OpenStudy (anonymous):

okay Thanks(: